# Learning on One Mode: Addressing Multi-Modality in Offline Reinforcement Learning

### v1

@@ 5 Practical Algorithm
17:     # Update the target networks.
@@ 5.1 Behaviour Policy Modelling
At the start of the LOM algorithm, we model the behaviour policy using a mixture density network (MDN) [Bishop, 1994]. The MDN receives a state $s$ and represents the resulting action distribution with a GMM, including its mixing coefficients, locations, and scales. We index the Gaussian components as the hyper-actions, ranging from $1$ to $M$.
@@ 5.3 Policy Learning
Finally, we learn the policy $\pi_{\theta}$ by optimising the objective:
$J(\theta)=\arg\max_{\pi}\mathbb{E}_{s\sim\mathcal{D},a\sim\pi_{\zeta_{g}}(\cdot\mid s)}\left[\exp_{clip}(\frac{1}{\beta}A^{\pi_{\zeta_{g}}}(s,a))\log\pi_{\theta}(a\mid s)\right],$ (18)
where $\beta$ is a hyper-parameter, $\exp_{clip}(\cdot)$ is the exponential function with a clipped range $(0,C]$, where $C$ is a positive number for numerical stability. We estimate the advantage using: $A^{\pi_{\zeta_{g}}}(s,a)\approx Q_{\psi}(s,a)-\mathbb{E}_{a\sim\pi_{\zeta_{g}}(\cdot\mid s)}[Q_{\psi}(s,a)]$, where we use the Q-function conditioned on the behaviour policy $\pi_{b}$ rather than $\pi_{\zeta_{g}}$ in order to reduce extrapolation error and computational complexity as well. Further, we employ the approach used in [Mao et al., 2023] to replace the expectation over $a\sim\pi_{\zeta_{g}}(\cdot\mid s)$ with the Q-value of the mean action of the Gaussian policy $\pi_{\zeta_{g}}$ to reduce computation.
@@ 6.1 LOM Achieves SOTA Performance in D4RL Benchmark
We evaluate LOM on three MuJoCo locomotion tasks from the D4RL benchmark [Fu et al., 2020]: halfcheetah, hopper, and walker2d. Each environment contains five dataset types: (i) medium — 1M samples from a policy trained to approximately one-third of expert performance; (ii) medium-replay — the replay buffer of a policy trained to match the performance of the medium agent (0.2M for halfcheetah, 0.4M for hopper, 0.3M for walker2d); (iii) medium-expert — a 50-50 split of medium and expert data (just under 2M samples); (iv) expert — 1M samples from a fully trained SAC policy [Haarnoja et al., 2018]; and (v) full-replay — 1M samples from the final replay buffer of an expert policy. Notably, the medium-replay and full-replay datasets are highly multi-modal, as states may have multiple action labels from different policies.
@@
Table 1: Averaged normalised scores on D4RL benchmarks over 4 random seeds. 00 represents the performance of a random policy and $100$ represents the performance of an expert policy. Standard deviations are provided as subscripts for multi-modal handling algorithms. The highest mean are highlighted in bold.
@@
Dataset Env OneStep AWAC CQL TD3BC LAPO DMPO WCGAN WCVAE LOM halfcheetah $50.4$ $47.9$ $47.0$ $48.3$ $45.9_{\pm 0.3}$ $47.5_{\pm 0.4}$ $48.2_{\pm 1.3}$ $50.5_{\pm 1.1}$ $\boldsymbol{51.0}_{\pm 0.7}$ medium hopper $87.5$ $59.8$ $53.0$ $59.3$ $51.6_{\pm 3.2}$ $71.2_{\pm 6.5}$ $78.6_{\pm 2.4}$ $89.0_{\pm 2.0}$ $\boldsymbol{100.8}_{\pm 1.4}$ walker2d $84.8$ $83.1$ $73.3$ $83.7$ $80.7_{\pm 0.8}$ $79.4_{\pm 4.7}$ $82.4_{\pm 1.7}$ $85.0_{\pm 1.5}$ $\boldsymbol{85.1}_{\pm 0.7}$ halfcheetah $42.7$ $44.8$ $45.5$ $44.6$ $44.7_{\pm 0.3}$ $45.2_{\pm 0.8}$ $42.3_{\pm 4.2}$ $45.0_{\pm 0.9}$ $\boldsymbol{48.8}_{\pm 0.7}$ medium-replay hopper $98.5$ $69.8$ $88.7$ $60.9$ $58.6_{\pm 3.8}$ $89.2_{\pm 8.1}$ $90.3_{\pm 3.1}$ $99.0_{\pm 2.1}$ $\boldsymbol{99.2}_{\pm 1.1}$ walker2d $61.7$ $78.1$ $81.8$ $81.8$ $71.7_{\pm 5.2}$ $82.1_{\pm 3.8}$ $72.6_{\pm 4.0}$ $74.0_{\pm 3.1}$ $\boldsymbol{84.8}_{\pm 1.0}$ halfcheetah $75.1$ $64.9$ $75.6$ $90.7$ $\boldsymbol{93.0}_{\pm 1.0}$ $91.1_{\pm 3.4}$ $76.5_{\pm 3.1}$ $80.0_{\pm 1.7}$ $92.7_{\pm 1.3}$ medium-expert hopper $108.6$ $100.1$ $105.6$ $98.0$ $105.2_{\pm 4.7}$ $78.4_{\pm 19.0}$ $110.0_{\pm 2.4}$ $109.0_{\pm 1.6}$ $\boldsymbol{110.1}_{\pm 1.4}$ walker2d $111.3$ $110.0$ $107.9$ $110.1_{\pm}$ $111.1_{\pm 0.2}$ $109.9_{\pm 0.4}$ $99.7_{\pm 1.0}$ $111.0_{\pm 0.7}$ $\boldsymbol{111.3}_{\pm 0.8}$ halfcheetah $88.2$ $81.7$ $96.3$ $96.7$ $95.9_{\pm 0.2}$ $\boldsymbol{97.0}_{\pm 1.0}$ $90.7_{\pm 1.7}$ $89.0_{\pm 0.8}$ $95.2_{\pm 0.4}$ expert hopper $106.9$ $109.5$ $96.5$ $107.8$ $106.7_{\pm 3.6}$ $93.6_{\pm 15.1}$ $107.3_{\pm 1.8}$ $108.0_{\pm 1.3}$ $\boldsymbol{111.0}_{\pm 0.8}$ walker2d $110.7$ $110.1$ $108.5$ $110.2$ $\boldsymbol{112.2}_{\pm 0.1}$ $111.4_{\pm 0.3}$ $109.3_{\pm 1.4}$ $111.0_{\pm 1.4}$ $109.3_{\pm 1.1}$ halfcheetah $64.7$ $66.6$ $73.6$ $71.7$ $74.2_{\pm 1.3}$ $71.4_{\pm 1.2}$ $64.2_{\pm 3.1}$ $66.0_{\pm 3.4}$ $\boldsymbol{76.6}_{\pm 1.2}$ full-replay hopper $69.8$ $100.1$ $98.2$ $76.8$ $100.0_{\pm 3.3}$ $101.5_{\pm 5.9}$ $80.2_{\pm 5.3}$ $86.5_{\pm 4.4}$ $\boldsymbol{102.0}_{\pm 2.7}$ walker2d $67.1$ $78.3$ $92.7$ $90.2$ $96.2_{\pm 2.8}$ $95.4_{\pm 1.8}$ $53.0_{\pm 5.4}$ $72.0_{\pm 3.7}$ $\boldsymbol{97.9}_{\pm 0.9}$
@@
Table 1 shows the benchmark results of our method against SOTA offline algorithms: OneStep [Brandfonbrener et al., 2021], AWAC [Nair et al., 2021], CQL [Kumar et al., 2020], and TD3BC [Fujimoto and Gu, 2021]. We also include four algorithms designed for handling multi-modality: LAPO [Chen et al., 2022], deterministic mixture policy optimisation (DMPO) [Osa et al., 2023], weighted conditioned GAN (WCGAN), and weighted conditioned VAE (WCVAE) [Wang et al., 2024b]. The experimental results are shown in Table 1, where the baselines’ results are mainly sourced from Mao et al. [2023] or reproduced using their official implementations. The results demonstrate that LOM outperforms the baseline algorithms in 12 out of 15 tasks. In tasks with multi-modal datasets, LOM surpasses all the baselines, and achieves a performance improvement ranging from $0.2\%$ to $7.9\%$ over the SOTA results.
@@ 6.2 Learning on One Mode Drives Improvements
We compare LOM with behaviour cloning (BC), AWAC, and MDN, introduced in Section 5.1. Specifically, BC and MDN imitate the behaviour policy using a unimodal Gaussian policy and a multi-modal Gaussian policy, respectively. AWAC applies weighted imitation learning to imitate the behaviour policy across all modes [Nair et al., 2021]. LOM, however, distinguishes multiple behaviour modes using MDN and focuses on learning from the best mode. Table 2 shows that LOM outperforms the baseline algorithms by an average of $53\%$.
@@ 7 Conclusions and Discussions
We observed that the hyper Q-function satisfies a Bellman-like equation, opening the possibility for learning it through Bellman updates. However, one challenge that emerged is the difficulty in accurately estimating the Bellman target for the hyper Q-function, as it requires computing the expectation over all possible hyper-actions, which may lead to extrapolation errors. Exploring a deterministic hyper-policy is a promising future direction to mitigate this issue, potentially simplifying the estimation process and further improving performance. Furthermore, the hyperparameter $M$ in LOM is intrinsically linked to the multi-modality present in the dataset. This parameter can be estimated using statistical techniques like bump hunting [Friedman and Fisher, 1999] and peak finding [Du et al., 2006], eliminating the need for fine-tuning. However, a key challenge lies in the fact that each state corresponds to a distinct multi-modal action distribution, making it computationally intensive to identify and count the modes for every individual state.
@@ Appendix A Motivating Example
To illustrate the challenge of capturing complex multi-modal action distributions, we present a toy example in Fig. 5. In this example, we generate a dataset where each state is associated with multiple valid actions, simulating the multi-modal nature of real-world offline RL datasets. We evaluate four models — Gaussian, Conditional Variational Auto-Encoder (CVAE) [Zhou et al., 2020, Chen et al., 2022], Conditional Generative Adversarial Networks (CGAN) [Yang et al., 2022b, Wang et al., 2024b], and Mixture Density Network (MDN) [Bishop, 1994] — on their ability to model the multi-modal action distribution.
@@ B.2 Proof of Theorem 1
Theorem 1. The composite policy induced by the greedy hyper-policy policy improves upon the behaviour policy:
@@ B.3 Proof of Theorem 2
As established in Wang et al. [2018], the following lemma provides a sufficient condition for proving that one policy is no worse than another. We restate this result as the following lemma:
Lemma (from Wang et al. [2018]): Suppose two policies $\pi$ and $\tilde{\pi}$ satisfy:
@@ B.4 Proof of Theorem 3
Following performance difference lemma [Kakade and Langford, 2002], we have
@@ C.1 Baselines
where $w(s,a)=\exp_{clip}\left(\frac{1}{\beta}A^{\pi_{b}}(s,a)\right)$. Here, $A^{\pi_{b}}$ represents the advantage of the behavior policy $\pi_{b}$. The WCGAN aims to generate action distributions that are weighted by this advantage, prioritising actions that yield higher returns.

### v2

@@ 5 Practical Algorithm
17:     # Update the target network.
@@ 5.1 Behaviour Policy Modelling
At the start of the LOM algorithm, we model the behaviour policy using a mixture density network (MDN) (Bishop, 1994). The MDN receives a state $s$ and represents the resulting action distribution with a GMM, including its mixing coefficients, locations, and scales. We index the Gaussian components as the hyper-actions, ranging from $1$ to $M$.
@@ 5.3 Policy Learning
Finally, we learn the policy $\pi_{\theta}$ by maximising the objective:
$J(\theta)=\mathbb{E}_{s\sim\mathcal{D},a\sim\pi_{\zeta_{g}}(\cdot\mid s)}\left[\exp_{clip}(\frac{1}{\beta}A^{\pi_{\zeta_{g}}}(s,a))\log\pi_{\theta}(a\mid s)\right],$ (18)
where $\beta$ is a hyper-parameter, $\exp_{clip}(\cdot)$ is the exponential function with a clipped range $(0,C]$, where $C$ is a positive number for numerical stability. We estimate the advantage using: $A^{\pi_{\zeta_{g}}}(s,a)\approx Q_{\psi}(s,a)-\mathbb{E}_{a\sim\pi_{\zeta_{g}}(\cdot\mid s)}[Q_{\psi}(s,a)]$, where we use the Q-function conditioned on the behaviour policy $\pi_{b}$ rather than $\pi_{\zeta_{g}}$ in order to reduce extrapolation error and computational complexity. Further, we employ the approach used in (Mao et al., 2023) to replace the expectation over $a\sim\pi_{\zeta_{g}}(\cdot\mid s)$ with the Q-value of the mean action of the Gaussian policy $\pi_{\zeta_{g}}$ to reduce computation.
@@ 6.1 LOM Achieves SOTA Performance in D4RL Benchmark
We evaluate LOM on three MuJoCo locomotion tasks from the D4RL benchmark (Fu et al., 2020): halfcheetah, hopper, and walker2d. Each environment contains five dataset types: (i) medium — 1M samples from a policy trained to approximately one-third of expert performance; (ii) medium-replay — the replay buffer of a policy trained to match the performance of the medium agent (0.2M for halfcheetah, 0.4M for hopper, 0.3M for walker2d); (iii) medium-expert — a 50-50 split of medium and expert data (just under 2M samples); (iv) expert — 1M samples from a fully trained SAC policy (Haarnoja et al., 2018); and (v) full-replay — 1M samples from the final replay buffer of an expert policy. Notably, the medium-replay and full-replay datasets are highly multi-modal, as states may have multiple action labels from different policies.
@@
Table 1: Averaged normalised scores on D4RL benchmarks over 4 random seeds. 00 represents the performance of a random policy and $100$ represents the performance of an expert policy. Standard deviations are provided as subscripts for multi-modal handling algorithms. The highest mean is highlighted in bold.
@@
Dataset Env OneStep AWAC CQL TD3BC LAPO DMPO WCGAN WCVAE LOM halfcheetah $50.4$ $47.9$ $47.0$ $48.3$ $45.9_{\pm 0.3}$ $47.5_{\pm 0.4}$ $48.2_{\pm 1.3}$ $50.5_{\pm 1.1}$ $\boldsymbol{51.0}_{\pm 0.7}$ medium hopper $87.5$ $59.8$ $53.0$ $59.3$ $51.6_{\pm 3.2}$ $71.2_{\pm 6.5}$ $78.6_{\pm 2.4}$ $89.0_{\pm 2.0}$ $\boldsymbol{100.8}_{\pm 1.4}$ walker2d $84.8$ $83.1$ $73.3$ $83.7$ $80.7_{\pm 0.8}$ $79.4_{\pm 4.7}$ $82.4_{\pm 1.7}$ $85.0_{\pm 1.5}$ $\boldsymbol{85.1}_{\pm 0.7}$ halfcheetah $42.7$ $44.8$ $45.5$ $44.6$ $44.7_{\pm 0.3}$ $45.2_{\pm 0.8}$ $42.3_{\pm 4.2}$ $45.0_{\pm 0.9}$ $\boldsymbol{48.8}_{\pm 0.7}$ medium-replay hopper $98.5$ $69.8$ $88.7$ $60.9$ $58.6_{\pm 3.8}$ $89.2_{\pm 8.1}$ $90.3_{\pm 3.1}$ $99.0_{\pm 2.1}$ $\boldsymbol{99.2}_{\pm 1.1}$ walker2d $61.7$ $78.1$ $81.8$ $81.8$ $71.7_{\pm 5.2}$ $82.1_{\pm 3.8}$ $72.6_{\pm 4.0}$ $74.0_{\pm 3.1}$ $\boldsymbol{84.8}_{\pm 1.0}$ halfcheetah $75.1$ $64.9$ $75.6$ $90.7$ $\boldsymbol{93.0}_{\pm 1.0}$ $91.1_{\pm 3.4}$ $76.5_{\pm 3.1}$ $80.0_{\pm 1.7}$ $92.7_{\pm 1.3}$ medium-expert hopper $108.6$ $100.1$ $105.6$ $98.0$ $105.2_{\pm 4.7}$ $78.4_{\pm 19.0}$ $110.0_{\pm 2.4}$ $109.0_{\pm 1.6}$ $\boldsymbol{110.1}_{\pm 1.4}$ walker2d $111.3$ $110.0$ $107.9$ $110.1$ $111.1_{\pm 0.2}$ $109.9_{\pm 0.4}$ $99.7_{\pm 1.0}$ $111.0_{\pm 0.7}$ $\boldsymbol{111.3}_{\pm 0.8}$ halfcheetah $88.2$ $81.7$ $96.3$ $96.7$ $95.9_{\pm 0.2}$ $\boldsymbol{97.0}_{\pm 1.0}$ $90.7_{\pm 1.7}$ $89.0_{\pm 0.8}$ $95.2_{\pm 0.4}$ expert hopper $106.9$ $109.5$ $96.5$ $107.8$ $106.7_{\pm 3.6}$ $93.6_{\pm 15.1}$ $107.3_{\pm 1.8}$ $108.0_{\pm 1.3}$ $\boldsymbol{111.0}_{\pm 0.8}$ walker2d $110.7$ $110.1$ $108.5$ $110.2$ $\boldsymbol{112.2}_{\pm 0.1}$ $111.4_{\pm 0.3}$ $109.3_{\pm 1.4}$ $111.0_{\pm 1.4}$ $109.3_{\pm 1.1}$ halfcheetah $64.7$ $66.6$ $73.6$ $71.7$ $74.2_{\pm 1.3}$ $71.4_{\pm 1.2}$ $64.2_{\pm 3.1}$ $66.0_{\pm 3.4}$ $\boldsymbol{76.6}_{\pm 1.2}$ full-replay hopper $69.8$ $100.1$ $98.2$ $76.8$ $100.0_{\pm 3.3}$ $101.5_{\pm 5.9}$ $80.2_{\pm 5.3}$ $86.5_{\pm 4.4}$ $\boldsymbol{102.0}_{\pm 2.7}$ walker2d $67.1$ $78.3$ $92.7$ $90.2$ $96.2_{\pm 2.8}$ $95.4_{\pm 1.8}$ $53.0_{\pm 5.4}$ $72.0_{\pm 3.7}$ $\boldsymbol{97.9}_{\pm 0.9}$
@@
Table 1 shows the benchmark results of our method against SOTA offline algorithms: OneStep (Brandfonbrener et al., 2021), AWAC (Nair et al., 2021), CQL (Kumar et al., 2020), and TD3BC (Fujimoto & Gu, 2021). We also include four algorithms designed for handling multi-modality: LAPO (Chen et al., 2022), deterministic mixture policy optimisation (DMPO) (Osa et al., 2023), weighted conditioned GAN (WCGAN), and weighted conditioned VAE (WCVAE) (Wang et al., 2024b). The experimental results are shown in Table 1, where the baselines’ results are mainly sourced from Mao et al. (2023) or reproduced using their official implementations. The results demonstrate that LOM outperforms the baseline algorithms in 12 out of 15 tasks. In tasks with multi-modal datasets, LOM surpasses all the baselines, and achieves a performance improvement ranging from $0.2\%$ to $7.9\%$ over the SOTA results.
@@ 6.2 Learning on One Mode Drives Improvements
We compare LOM with behaviour cloning (BC), AWAC, and MDN, introduced in Section 5.1. Specifically, BC and MDN imitate the behaviour policy using a unimodal Gaussian policy and a multi-modal Gaussian policy, respectively. AWAC applies weighted imitation learning to imitate the behaviour policy across all modes (Nair et al., 2021). LOM, however, distinguishes multiple behaviour modes using MDN and focuses on learning from the best mode. Table 2 shows that LOM outperforms the baseline algorithms by an average of $53\%$.
@@ 7 Conclusions and Discussions
We observed that the hyper Q-function satisfies a Bellman-like equation, opening the possibility for learning it through Bellman updates. However, one challenge that emerges is the difficulty in accurately estimating the Bellman target for the hyper Q-function, as it requires computing the expectation over all possible hyper-actions, which may lead to extrapolation errors. Exploring a deterministic hyper-policy is a promising future direction to mitigate this issue, potentially simplifying the estimation process and further improving performance. Furthermore, the hyperparameter $M$ in LOM is intrinsically linked to the multi-modality present in the dataset. This parameter can be estimated using statistical techniques like bump hunting (Friedman & Fisher, 1999) and peak finding (Du et al., 2006), eliminating the need for fine-tuning. However, a key challenge lies in the fact that each state corresponds to a distinct multi-modal action distribution, making it computationally intensive to identify and count the modes for every individual state.
@@ Appendix A Motivating Example
To illustrate the challenge of capturing complex multi-modal action distributions, we present a toy example in Fig. 5. In this example, we generate a dataset where each state is associated with multiple valid actions, simulating the multi-modal nature of real-world offline RL datasets. We evaluate four models — Gaussian, Conditional Variational Auto-Encoder (CVAE) (Zhou et al., 2020; Chen et al., 2022), Conditional Generative Adversarial Networks (CGAN) (Yang et al., 2022b; Wang et al., 2024b), and Mixture Density Network (MDN) (Bishop, 1994) — on their ability to model the multi-modal action distribution.
@@ B.2 Proof of Theorem 1
Theorem 1. The composite policy induced by the greedy hyper-policy improves upon the behaviour policy:
@@ B.3 Proof of Theorem 2
As established in Wang et al. (2018), the following lemma provides a sufficient condition for proving that one policy is no worse than another. We restate this result as the following lemma:
Lemma (from Wang et al. (2018)): Suppose two policies $\pi$ and $\tilde{\pi}$ satisfy:
@@ B.4 Proof of Theorem 3
Following performance difference lemma (Kakade & Langford, 2002), we have
@@ C.1 Baselines
where $w(s,a)=\exp_{clip}\left(\frac{1}{\beta}A^{\pi_{b}}(s,a)\right)$. Here, $A^{\pi_{b}}$ represents the advantage of the behaviour policy $\pi_{b}$. The WCGAN aims to generate action distributions that are weighted by this advantage, prioritising actions that yield higher returns.